OpenStudy (anonymous):

help please? what is 2.06666666667 written as a fraction

OpenStudy (anonymous):

2 and a sixteenth. This could be written as: 2+(1/16) OR 33/16.

OpenStudy (andriod09):

simple:\[\frac {2.07}{10}\]

OpenStudy (anonymous):

The reason being 0.6666666... is a sixteenth, and two is equal to 32 sixteenths.

OpenStudy (anonymous):

Please rate mine as the top answer! :)

OpenStudy (raden):

well. i want to try explain to u let : x = 2.066666... (1) if u multiply by 10 to both sides, giving us 10x = 20.66666... (2) and if u multiply by 100 to both sides (the first equation), it can be 100x = 206.66666... (3) now, subtract the (3) - (2) eq 100x-10x = 206.66666... - 20.66666... 90x = 186 x = 186/90 x = 31/15

OpenStudy (anonymous):

You can also figure it out using infinite series, but that takes a bit longer.

OpenStudy (phi):

@Dthnider 0.666... is 2/3, and 0.066666... is 2/3 * 1/10 or 2/30 = 1/15 the answer is 2 and 1/15 see RadEn's post on how to solve these types of problems